ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 162, 2007

Abért, M. and P. Hegedűs On the number of generators needed for free profinite products of finite groups

81

Brešar, M. Commutativity preserving maps revisited  317
Deng, S. and Z. Hou On symmetric Finsler spaces  197

Deville, R. and P. Hájek Smooth noncompact operators from C(K), K scattered

29

Fernández-Alcober, G. A. Omega subgroups of powerful p-groups  75

De Filippis, V. An Engel condition with generalized derivations on multilinear polynamials

93

Hájek P. See Deville, R.

Harbater, D. Correction and addendum to “Embedding problems with local conditions”

373

Hegedűs, P. See M. Abért

Hou, Z. See S. Deng

Irmak, E. and M. Korkmaz Automorphisms of the Hatcher-Thurston complex

183

Kalenda, O. F. K. (I)-envelopes of closed convex sets in Banach spaces 157
Kalton, N. J. Extending Lipschitz maps into C(K)-spaces  275
Kollár, J. A conjecture of Ax and degenerations of Fano varieties  235

Korkmaz, M. See E. Irmak

Li, B. Qin and E. G. Saleeby On solutions of functional-differential equations f′(x) = a(x)f(g(x)) + b(x)f(x) + c(x) in the large

335

Li, Y.-C., R. Sato and S.-Y. Shaw Convergence theorems and Tauberian theorems for functions and sequences in Banach spaces and Banach lattices

109

Natale, S. On the exponent of tensor categories coming from finite groups

253

Rosendal, C. and S. Solecki Automatic continuity of homomorphisms and fixed points on metric compacta 349 Saleeby, E. G. See B. Qin Li Sato, R. See Y.-C. Li Shaw, S.-Y. See Y.-C. Li Solecki, S. See C. Rosendal

Spinas, O. Splitting squares  57
de Thélin, H. Construction of measures with expansion  151

Wang, X. Algebraic convergence theorems of n-dimensional Kleinian groups

221

Yekutieli, A. Mixed resolutions and simplicial sections  1